OpenStudy (anonymous):

how do i find the properties of this hyperbola -4x^2+y^2-8x-12y+36=0

OpenStudy (amistre64):

generally by completing the squares for the x and y parts

OpenStudy (anonymous):

can you show me how to work the problem out

OpenStudy (amistre64):

yes, but it requires knowledge of completing a square; do you remember how to work that?

OpenStudy (anonymous):

yes little bit i would have to group my common factors first right?

OpenStudy (amistre64):

common x and y parts yes -4x^2+y^2-8x-12y+36=0 (-4x^2-8x + ___ ) + (y^2-12y + ___ ) +36 = 0

OpenStudy (amistre64):

looks like the 36 already completes the y square for you

OpenStudy (amistre64):

(-4x^2-8x + ___ ) + (y^2-12y + 36) = 0 (-4x^2-8x + ___ ) + (y-6)^2 = 0

OpenStudy (anonymous):

-24?

OpenStudy (amistre64):

dunno, lets factor out that -4 from the x parts (-4x^2-8x + ___ ) + (y-6)^2 = 0 -4 (x^2+2x + ___ ) + (y-6)^2 = 0 id say +1 -1 -4 (x^2+2x + 1 - 1) + (y-6)^2 = 0 -4 (x^2+2x + 1) +4 + (y-6)^2 = 0 -4 (x+1)^2 + (y-6)^2 = -4

OpenStudy (amistre64):

divide both sides by -4 and we have ... (x+1)^2 (y-6)^2 ------- - ------- = 1 1 4

OpenStudy (anonymous):

now its starting to come back to me

OpenStudy (anonymous):

after this step i get lost

OpenStudy (amistre64):

well, we can determine alot of properties from this now; it all depends what properties you need to find

OpenStudy (anonymous):

foci

OpenStudy (amistre64):

well, we would need to determine which way this opens up; since -y^2 can never be positive, it opens along the (+-x,0) parts

OpenStudy (amistre64):

lets go ahead and view this from the origin by taking out the center bits x^2 y^2 --- - ---- = 1 1^2 2^2 when y=0, x = -1, +1  so something like this